OpenStudy (anonymous):

7x+5y=7 over 2 y=2x-1 What is the ordered pair?

OpenStudy (anonymous):



OpenStudy (ranga):

First put the equations in the right order: 7x + 5y = 7/2 -2x + y = -1 Multiply second equation by -5 and add it to the first. That will get rid of the y term and you can solve for x. Then put the x value in the second equation and solve for y.